OpenStudy (anonymous):

I still do not understand how to do this: Find the zeros of the function: y=(x-2)(x+9)... Help....... :/

OpenStudy (anonymous):

well the thing is someone already factor a quadratic for you. (x-2)(x+9)=0 results in x-2=0 and x+9=0 solve each individually. x=2 and x=-9

OpenStudy (anonymous):

so the -2 and the 9 just switch signs?

OpenStudy (anonymous):

yes. I just solve for them individuallay x-2=0 add 2 to both sides to isolate the x x+9=0 subtract 9 to both sides to isolate the x ------- here is the graph https://www.google.com/search?q=(x-2)(x%2B9)&rlz=1C1CHFX_enUS470US470&aq=f&sugexp=chrome,mod=11&sourceid=chrome&ie=UTF-8 Notice it intercepts the x axis the points i solved for

OpenStudy (anonymous):

oh yes okay i see that

OpenStudy (anonymous):

yay :D

OpenStudy (anonymous):

Haha Thank you (:

OpenStudy (anonymous):

de nada